OpenStudy (anonymous):

what is the difference between cross and dot product? can the dot product be used to determine if vectors are collinear?

OpenStudy (anonymous):

let a and b be two vectors making an angle theta with each other. \[then a \times b is called cross product and a.b is called \dot product.\]

OpenStudy (anonymous):

\[a \times b=\left| a \right|\left| b \right|\sin \theta n,where n ia unit vector perpendicular to both a and b\]

OpenStudy (anonymous):

\[a.b =\left| a \right|\left| b \right|\cos \theta \]

OpenStudy (anonymous):

thanks in what way is n multiplied to the cross formula? simple multiplication?

OpenStudy (anonymous):

when the vectors are collinear angle between them is 0 or pi. cos 0=1,and cos pi=-1 so it can be used as well.

OpenStudy (anonymous):

cross product is a vector quantity.

OpenStudy (anonymous):

I see. in contrast, dot product is a scalar quantity

OpenStudy (anonymous):

yea